# MagicDrive3D: Controllable 3D Generation for Any-View Rendering in Street Scenes

### v3

@@ Abstract
While controllable generative models for images and videos have achieved remarkable success, high-quality models for 3D scenes, particularly in unbounded scenarios like autonomous driving, remain underdeveloped due to high data acquisition costs. In this paper, we introduce MagicDrive3D, a novel pipeline for controllable 3D street scene generation that supports multi-condition control, including BEV maps, 3D objects, and text descriptions. Unlike previous methods that reconstruct before training the generative models, MagicDrive3D first trains a video generation model and then reconstructs from the generated data. This innovative approach enables easily controllable generation and static scene acquisition, resulting in high-quality scene reconstruction. To address the minor errors in generated content, we propose deformable Gaussian splatting with monocular depth initialization and appearance modeling to manage exposure discrepancies across viewpoints. Validated on the nuScenes dataset, MagicDrive3D generates diverse, high-quality 3D driving scenes that support any-view rendering and enhance downstream tasks like BEV segmentation. Our results demonstrate the framework’s superior performance, showcasing its potential for autonomous driving simulation and beyond.
@@ 1 Introduction
With the advancement of generative models [10, 11, 25, 22], there has been increasing interest in generating 3D assets [19, 28, 2]. While many works focus on object-centric generation [19, 27], generating open-ended 3D scenes remains relatively unexplored. This gap is even more critical considering many downstream applications, such as virtual reality and autonomous driving, require controllable generation of 3D street scenes, which is an open challenge.
The typical workflow for 3D scene generation is geometry-focused [^1] In this paper, we focus on generative models where views/scenes are generated from latent variables.. It begins with reconstructing the scene from multi-view sensors to obtain a compact representation (e.g., voxel grids in NF-LDM [13] and tri-planes in GAUDI [2]), followed by training generative models to produce the scene representation from latent variables. Despite their broad applicability, these methods require expensive data collection for the initial reconstruction step, which sets the upper limit on generation quality. Specifically, they often require static scenes and consistent sensor properties, such as exposure and white balance. Street view datasets, like nuScenes [4], fail to meet these requirements, making it extremely challenging to train geometry-focused 3D street scene generation models using such datasets.
Geometry-free methods [21] represent another category in 3D-aware view synthesis. In street-view generation, multi-view 2D images or videos can be directly generated based on camera parameters [9, 32, 30], excelling in photorealistic image generation. However, they often lack sufficient geometric consistency, limiting their ability to extend to viewpoints beyond the dataset [9, 32, 30].
Recognizing the advancements in controllable generation by geometry-free view synthesis methods [5, 9], it is potential to use them as data engines. Their controllability and photo-realism could address the challenges faced by geometry-focused methods. However, the limited 3D consistency in synthetic views from geometry-free methods, such as temporal inconsistency among frames and deformation of objects, poses crucial issues for integrating two kinds of methods into a unified framework.
@@
To address these challenges, we propose MagicDrive3D, a novel framework that combines geometry-free view synthesis and geometry-focused reconstruction for controllable 3D street scene generation. As illustrated in Figure 2, our approach begins with training a multi-view video generation model to synthesize multiple views of a static scene. This model is configured using controls from object boxes, road maps, text prompts, and camera poses. To enhance inter-frame 3D consistency, we incorporate coordinate embeddings that represent the relative transformation between LiDAR coordinates for accurate control of frame positions. Next, we improve the reconstruction quality of generated views by enhancing 3D Gaussian splatting from the perspectives of prior knowledge, modeling, and loss functions. Given the limited overlap between different camera views [34], we adopt a monocular depth prior and propose a dedicated algorithm for alignment in sparse-view settings. Additionally, we introduce deformable Gaussian splatting and appearance embedding maps to handle local dynamics and exposure discrepancies, respectively.
@@
Demonstrated by extensive experiments, our MagicDrive3D framework excels in generating highly realistic street scenes that align with road maps, 3D bounding boxes, ego trajectory, and text descriptions, as exemplified in Figure 1. We show that the generated camera views can augment training for Bird’s Eye View (BEV) segmentation tasks, improving the view-point robustness on perception tasks. Notably, MagicDrive3D is the first to achieve controllable 3D street scene generation using a common driving dataset (e.g., the nuScenes dataset [4]), without requiring repeated data collection from static scenes.
@@
  • We propose MagicDrive3D, the first framework to effectively integrate both geometry-free and geometry-focused view synthesis for controllable 3D street scene generation. MagicDrive3D generates realistic 3D street scenes that support rendering from any camera view according to various control signals.
  • We introduce a relative pose embedding technique to generate videos with improved 3D consistency. Additionally, we enhance the reconstruction quality with tailored techniques, including deformable Gaussian splatting, to handle local dynamics and exposure discrepancies in the generated videos.
  • Through extensive experiments, we demonstrate that MagicDrive3D generates high-quality street scenes with multi-dimensional controllability. Our results also show that synthetic data improves 3D perception tasks, highlighting the practical benefits of our approach.
@@ 2 Related Work
3D Scene Generation. Numerous 3D-aware generative models can synthesize images with explicit camera pose control [38, 21] and potentially other scene properties [26], but only a few scale for open-ended 3D scene generation. GSN [7] and GAUDI [2], representative of models generating indoor scenes, utilize NeRF [18] with latent code input for “floorplan” or tri-plane feature. Their reliance on datasets covering different camera poses is incompatible with typical driving datasets where camera configuration remains constant. NF-LDM [13] develops a hierarchical latent diffusion model for scene feature voxel grid generation. However, their representation and complex modeling hinder fine detail generation.
Contrary to previous works focusing on scene generation using explicit geometry, often requiring substantial data not suitable for typical street view datasets (e.g., nuScenes [4]) as discussed in Section 1, we propose merging geometry-free view synthesis with geometry-focused scene representations for controllable street scene creation. Methodologically, LucidDreamer [6] is most similar to our approach, although it relies on a text-controlled image generation model, which cannot qualify as a view synthesis model. In contrast, our video generation model is 3D-aware. Besides, we propose several improvements over 3DGS for better scene generation quality. Besides, inpainting-based methods like LucidDreamer [6] and WonderJourney [36] cannot complete our controllable street scene generation task. We showcase the differences in Appendix E.
Street View Video Generation. Diffusion models [25, 11] boost a series of works for street view video generation. These works span from single view video (e.g., [29]) to multi-view video (e.g., [9, 32, 30]). Despite cross-view consistency being essential for multi-view video generation, their generalization ability on camera poses is limited due to the data-centric nature [9]. Furthermore, these models lack exact control over frame transformation (i.e., precise ego trajectory), which is crucial for scene reconstruction. Our work addresses this by enhancing control in video generation and proposing a dedicated deformable Gaussian splatting for geometric assurance.
Street Scene Reconstruction. Scene reconstruction and novel view rendering for street views are useful in applications like driving simulation, data generation, and augmented and virtual reality. For street scenes, attributes like scene dynamic and discrepancies from multi-camera data collection make typical large-scale reconstruction methods ineffective (e.g., [20, 16, 15]). Hence, real data-based reconstruction methods like [34, 35] utilize LiDAR for depth prior, but their output only permits novel view rendering from the same scene. Unlike these methods, our approach synthesizes novel scenes under multiple levels of conditional controls.
@@ 3 Preliminaries
Problem Formulation. In this paper, we focus on controllable street scene generation. Given scene description $\mathbf{S}_{t}$, our task is to generate street scenes (represented with 3D Gaussians $\mathbf{G}$) that correspond to the description from a set of latent $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e. $\mathbf{G}=\mathcal{G}(\mathbf{S}_{t},\mathbf{z})$. To describe a street scene, we adopt the most commonly used conditions as per [9, 30, 32]. Specifically, a frame of driving scene $\mathbf{S}_{t}=\{\mathbf{M}_{t},\mathbf{B}_{t},\mathbf{L}_{t}\}$ is described by road map $\mathbf{M}_{t}\in\{0,1\}^{w\times h\times c}$ (a binary map representing a $w\times h$ meter road area in BEV with $c$ semantic classes), 3D bounding boxes $\mathbf{B}_{t}=\{(c_{i},b_{i})\}_{i=1}^{N}$ (each object is described by box $b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$ and class $c_{i}\in\mathcal{C}$), and text $\mathbf{L}_{t}$ describing additional information about the scene (e.g., weather and time of day). In this paper, we parameterize all geometric information according to the LiDAR coordinate of the ego car.
One direct application of scene generation is any-view rendering. Specifically, given any camera pose $\mathbf{P}=[\mathbf{K},\mathbf{R},\mathbf{t}]$ (i.e., intrinsics, rotation, and translation), the model $\mathbf{G}(\cdot)$ should render photo-realistic views with 3D consistency, $\mathcal{I}^{r}=\mathbf{G}(\mathbf{P})$, which is not applicable to previous street view generation (e.g., [9, 30, 32]). Besides, we present more applications in Section 5.
@@ 4 Methods
In this section, we introduce our controllable street scene generation pipeline. Due to the challenges that exist in data collection, we integrate geometry-free view synthesis and geometry-focused reconstruction, and propose a generation-reconstruction pipeline, detailed in Section 4.1 and Figure 2. Specifically, we introduce a controllable video generation model to connect control signals with camera views (Section 4.2) and enhance the 3DGS from prior, modeling and loss perspectives (Section 4.3) for better reconstruction with generated views.
Figure 2: Method Overview of MagicDrive3D. For controllable street scene generation, MagicDrive3D decomposes the task into two steps: ① conditional multi-view video generation, which tackles the control signals and provides detailed prior of the scene; and ② scene reconstruction with deformable Gaussian splatting, which guarantees view consistency for any-view rendering.
4.1 3D Street Scene Generation
@@
Direct modeling of controllable street scene generation faces two major challenges: scene dynamics and discrepancy in data collection. Scene dynamics refer to the movements and deformation of elements in the scene, while discrepancy in data collection refer to the discrepancy (e.g., exposure) caused by data collection. These two challenges are even more severe due to the sparsity of cameras for street views (e.g., typically only 6 surrounding perspective cameras). Therefore, reconstruction-generation frameworks do not work well for street scene generation [13, 2].
@@
Figure 2 shows the overview of MagicDrive3D. Given scene descriptions $\mathbf{S}$ as input, MagicDrive3D first extend the descriptions into sequence $\{\mathbf{S}_{t}\}$, where $t\in[0,T]$ according to preset camera poses $\{\mathbf{P}_{c,t}\}$, and generate a sequence of successive multi-view images $\{\mathcal{I}_{c,t}\}$, where $c\in\{1,\dots,N\}$ refers to $N$ surrounding cameras, according to conditions $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$ (detailed in Section 4.2). Then we construct Gaussian representation of the scene with $\{\mathcal{I}_{c,t}\}$ and camera poses $\{\mathbf{P}_{c,t}\}$ as input. This step contains an initializing procedure with a pre-trained monocular depth model and an optimizing process with deformable Gaussian splatting (detailed in Section 4.3). Consequently, the generated street scene not only supports any-view rendering, but also accurately reflects different control signals.
@@
MagicDrive3D integrates geometry-free view synthesis and geometry-focused reconstruction, where control signals are tackled by a multi-view video generator, while the reconstruction step guarantees the generalization ability for any-view rendering. Such a video generator has two advantages: first, since multi-view video generation does not require generalization on novel views [9], it poses less data dependency for street scenes; Second, through conditional training, the model is able to decompose control signals, turning dynamic scenes into static scenes that are more friendly for reconstruction. Besides, for the reconstruction step, strong prior from the multi-view video reduces the burden for scene modeling with complex details.
4.2 Relative Pose Control for Video Generation

### v4

@@ Abstract
Controllable generative models for images and videos have seen significant success, yet 3D scene generation, especially in unbounded scenarios like autonomous driving, remains underdeveloped. Existing methods lack flexible controllability and often rely on dense view data collection in controlled environments, limiting their generalizability across common datasets (e.g., nuScenes). In this paper, we introduce MagicDrive3D, a novel framework for controllable 3D street scene generation that combines video-based view synthesis with 3D representation (3DGS) generation. It supports multi-condition control, including road maps, 3D objects, and text descriptions. Unlike previous approaches that require 3D representation before training, MagicDrive3D first trains a multi-view video generation model to synthesize diverse street views. This method utilizes routinely collected autonomous driving data, reducing data acquisition challenges and enriching 3D scene generation. In the 3DGS generation step, we introduce Fault-Tolerant Gaussian Splatting to address minor errors and use monocular depth for better initialization, alongside appearance modeling to manage exposure discrepancies across viewpoints. Experiments show that MagicDrive3D generates diverse, high-quality 3D driving scenes, supports any-view rendering, and enhances downstream tasks like BEV segmentation, demonstrating its potential for autonomous driving simulation and beyond.
@@ 1 Introduction
The advancement of generative models has enabled their use as data engines, particularly in the realm of controllable generative models, which can produce synthetic data without the need for relabeling [5]. This capability is especially valuable in downstream applications like autonomous driving. However, much of the current research is limited to the generation of 2D data, such as images [5, 33] and videos [11, 10], which falls short of meeting the demands for data that requires 3D information, such as viewpoint changes [15] or post-generation 3D editing [6]. Consequently, controllable generation of 3D street scenes remains an open challenge.
Table 1: Comparisons on data collection requirements and functionality among different types of methods. MagicDrive3D belongs to 3D scene generation but can do more and requires less than previous works. Due to space limits, we only cite some typical works for each category. More detailed related works can be found in Section 2.
Method Type Data Collection Requirements
Prior
Control
Novel
Appearance
Novel Pose Rendering
Allow
Dynamic Scene
Allow Sparse
Pose Coverage
Within Trajectory 360
Reconstruction [18, 30, 40, 35, 44, 22, 48, 6]
Image/Video Generation [11, 10]
3D Scene Generation ([14, 2])
3D Scene Generation (MagicDrive3D)
3D generation [^1] In this paper, we focus on generative models where views/scenes are generated from latent variables, which is different from reconstruction (e.g., [38]) or view synthesis (e.g., [25])., on the other hand, can directly produce 3D assets based on given conditions [29, 41], supporting data engine applications that require 3D information. However, many of them focus on single objects [29] or creative scenes [24] for gaming or content creation. Only a few attempts have been made at realistic scene generation for the physical world [14, 2]; for example, NF-LDM [14] enables controllable synthesis of 3D scenes for autonomous driving. The typical workflow for 3D generation involves first obtaining a compact 3D representation (e.g., voxel grids in [14]) and then training a generative model based on this representation. Due to the high demands for data acquisition and reconstruction quality (e.g., the need for static scenes and numerous overlapping camera views), these methods are often limited to specific autonomous driving scenarios (e.g., parking lots, where no objects are moving), and thus do not support common street view datasets (e.g., nuScenes [4]).
These limitations prompt us to reconsider the relationship between the typical 3D generation workflow and 2D controllable generative models in autonomous driving scene generation. We observe that: 1) 2D controllable generation requires less on data acquisition, as common roadside data can support training; 2) 2D controllable generation provides the ability to manipulate data (e.g., generating multiple viewpoints of static scenes) to meet the data requirements of 3D reconstruction algorithms; 3) 3D representation is the core of 3D generation to achieve consistent view rendering and novel view generalization. These observations inspire us to restructure the 3D scene generation workflow by using 2D controllable generation as the initial step, leveraging its data manipulation capabilities to create reconstruction-friendly camera view data, and finally obtain the 3D representation of the scene. This ensures consistent viewpoint rendering and supports applications requiring 3D information (e.g., viewpoint augmentation [15]).
@@
In this paper, we propose MagicDrive3D, a novel framework that combines video-based view synthesis and 3D representation (3DGS) generation for controllable 3D street scene generation, as illustrated in Figure 1. Our approach begins with training a multi-view video generation model to synthesize multiple street views. Besides text prompts, this model is configured using conditions from object boxes, road maps, and camera poses to support precise content control for static scene transformation. Additionally, we incorporate ego trajectory embeddings representing the relative pose transformation between each frame, eliminating the need for pose estimation after view generation, leading to reconstruction-friendly view generation. Next, we identify and solve several drawbacks in existing reconstruction algorithms on generated views. Since artifacts in view generation break the consistency assumption of reconstruction algorithms (e.g., 3DGS [13]), these artifacts can be amplified in the generated 3D representation, resulting in degraded rendering quality. To solve this issue, we improve the 3D representation generation from the generated views from the perspectives of prior knowledge, modeling, and loss functions. Most importantly, we introduce Fault-Tolerant Gaussian Splatting (FTGS) and appearance embedding maps to handle local inconsistency and exposure discrepancies, respectively.
@@
Demonstrated by extensive experiments, our MagicDrive3D framework excels in generating highly realistic street scenes that align with road maps, 3D bounding boxes, and text descriptions, as exemplified in Figure 1. We further show that the generated camera views can augment training of Bird’s Eye View (BEV) segmentation tasks, improving the viewpoint robustness on perception tasks. Notably, MagicDrive3D is the first to achieve controllable 3D street scene generation using a common driving dataset (e.g., the nuScenes dataset [4]), without requiring repeated data collection from static scenes or other sensor information.
@@
  • We propose MagicDrive3D, the first framework to effectively integrate both view synthesis and 3D representation generation for controllable 3D street scene generation. MagicDrive3D generates realistic 3DGS for street scenes that support rendering from any camera view according to various control signals.
  • We introduce a relative pose embedding to generate videos with accurate pose controls. Besides, we enhance the 3D representation generation quality through a dedicated pipeline (including Fault-Tolerant Gaussian Splatting) to handle local inconsistency and exposure discrepancies in the generated videos.
  • Through extensive experiments, we demonstrate that MagicDrive3D generates high-quality street scenes with various controllability. We also show that synthetic data improves 3D perception tasks, highlighting the practical benefits of our approach.
@@ 2 Related Work
3D Scene Generation. Numerous 3D-aware generative models can synthesize images with explicit camera pose control [45, 25] and potentially other scene properties [29], but only a few scale for open-ended 3D scene generation. GSN [8] and GAUDI [2], representative of models generating indoor scenes, utilize NeRF [21] with latent code input for “floorplan” or tri-plane feature. Their reliance on datasets covering different camera poses is incompatible with typical driving datasets where only fixed cameras are available. NF-LDM [14] develops a hierarchical latent diffusion model for scene feature voxel grid generation. However, their representation and complex modeling hinder fine detail generation.
Contrary to previous works focusing on scene generation using explicit geometry, often requiring substantial data not suitable for typical street view datasets (e.g., nuScenes [4]) as discussed in Section 1, we propose merging geometry-free view synthesis with geometry-focused scene representations for controllable street scene creation. Table 1 shows a detailed comparison among different types of methods. Methodologically, inpainting-based methods [7, 16, 42] are most similar to our approach. However, they generally rely on text-controlled image generation models, which cannot qualify as a view synthesis model. Besides, they cannot complete our street scene generation task due to the lack of controllability, as showcased in Appendix G. In contrast, our video generation model is 3D-aware. Besides, we propose several improvements over 3DGS for better scene generation quality.
2D Street View Generation. Diffusion models [28, 12] boost a series of works for 2D street view generation. These works span from single view video (e.g., [31]) to multi-view video (e.g., [11, 36, 32]). Despite cross-view consistency being essential for multi-view video generation, their generalization ability on camera poses is limited due to the data-centric nature [11]. Furthermore, these models lack accurate control over frame transformation (e.g., precise ego trajectory), which is crucial for 3DGS generation. Our work addresses this by enhancing pose control in video generation and proposing a dedicated 3DGS generation pipeline for geometric assurance.
Street Scene Reconstruction. Scene reconstruction (or novel view rendering) for street views is useful in applications like driving simulation, and augmented and virtual reality. For street scenes, attributes like scene dynamic and discrepancies from multi-camera data collection make typical large-scale reconstruction methods ineffective (e.g., Rematas et al. [23], Martin-Brualla et al. [19], Lin et al. [17]). Hence, real data-based reconstruction methods like Xie et al. [38], Yan et al. [39] utilize LiDAR for depth prior, but their output only permits novel view rendering, lacking the capability of novel scene generation. Unlike these methods, our approach synthesizes novel scenes under multiple levels of conditional controls.
@@ 3 Preliminaries
Problem Formulation. In this paper, we focus on controllable street scene generation. Given scene description $\mathbf{S}_{t}$, our task is to generate street scenes (represented with 3D Gaussians $\mathbf{G}$) that correspond to the description from a set of latent $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e. $\mathbf{G}=\mathcal{G}(\mathbf{S}_{t},\mathbf{z})$. To describe a street scene, we adopt the most commonly used conditions as per Gao et al. [11], Wang et al. [32], Wen et al. [36]. Specifically, a frame of driving scene $\mathbf{S}_{t}=\{\mathbf{M}_{t},\mathbf{B}_{t},\mathbf{L}_{t}\}$ is described by road map $\mathbf{M}_{t}\in\{0,1\}^{w\times h\times c}$ (a binary map representing a $w\times h$ meter road area in BEV with $c$ semantic classes), 3D bounding boxes $\mathbf{B}_{t}=\{(c_{i},b_{i})\}_{i=1}^{N}$ (each object is described by box $b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$ and class $c_{i}\in\mathcal{C}$), and text $\mathbf{L}_{t}$ describing additional information about the scene (e.g., weather and time of day). In this paper, we parameterize all geometric information according to the LiDAR coordinate of the ego car.
One direct application of scene generation is any-view rendering. Specifically, given any camera pose $\mathbf{P}=[\mathbf{K},\mathbf{R},\mathbf{t}]$ (i.e., intrinsics, rotation, and translation), the 3DGS model $\mathbf{G}(\cdot)$ should render photo-realistic views of the generated scene with 3D consistency, $\mathcal{I}^{r}=\mathbf{G}(\mathbf{P})$, which is not applicable to previous street view generation (e.g., Gao et al. [11], Wang et al. [32], Wen et al. [36]). Besides, we present more practical applications in Section 5.
Figure 2: Method Overview of MagicDrive3D. For controllable street 3D scene generation, MagicDrive3D decomposes the task into two steps: ① conditional multi-view video generation, which tackles the control signals and generates consistent view priors to the novel scene; and ② Gaussian Splatting generation with our Enhanced GS pipeline, which supports various viewpoint rendering (e.g., panorama).
@@ 4 Methods
In this section, we introduce our controllable street scene generation pipeline. Due to the challenges that exist in data collection, we integrate geometry-free view synthesis and geometry-focused 3DGS generation, detailed in Section 4.1 and Figure 2. Specifically, we introduce a controllable video generation model to connect control signals and 3DGS generation by camera views generation (Section 4.2). Besides, we present a dedicated 3DGS generation pipeline for generated views, enhancing from prior, modeling, and loss perspectives (Section 4.3).
4.1 Overview of MagicDrive3D
@@
Direct modeling of controllable street scene generation faces two major challenges: scene dynamics and discrepancy in data collection. Scene dynamics refers to the movements and deformation of elements in the scene, while discrepancy in data collection refers to the discrepancy (e.g., exposure) caused by data collection. These two challenges are even more severe due to the sparsity of cameras for street views (e.g., typically only 6 surrounding perspective cameras). Therefore, reconstruction-generation frameworks do not work well for street scene generation [14, 2].
@@
Figure 2 shows the overview of MagicDrive3D. Given scene descriptions $\mathbf{S}$ as input, MagicDrive3D first extend the descriptions into sequence $\{\mathbf{S}_{t}\}$, where $t\in[0,T]$ according to preset camera poses $\{\mathbf{P}_{c,t}\}$, and generate a sequence of successive multi-view images $\{\mathcal{I}_{c,t}\}$, where $c\in\{1,\dots,N\}$ refers to $N$ surrounding cameras, according to conditions $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$ (detailed in Section 4.2). Then we generate 3D Gaussian representation of the scene with $\{\mathcal{I}_{c,t}\}$ and camera poses $\{\mathbf{P}_{c,t}\}$ as input. This step contains an initializing procedure with a pre-trained monocular depth model and an optimizing process with Fault-Tolerant Gaussian Splatting (FTGS, detailed in Section 4.3). Consequently, the generated street scene reflects different control signals accurately and supports any-view rendering.
@@
MagicDrive3D integrates geometry-free view synthesis and GS-based scene representation, which are geometry-focused. Specifically, control signals are tackled by a multi-view video generator, while the 3DGS generation step guarantees the generalization ability for any-view rendering. Such a video generator has two advantages: first, since multi-view video generation does not require generalization on novel views [11], it poses less data dependency for street scenes; second, through conditional training, the model is able to decompose control signals, turning dynamic scenes into static scenes that are more friendly for 3DGS generation. Besides, for the 3DGS generation step, strong priors from the multi-view video reduce the burden for scene modeling with complex details.
Figure 3: Illustration of the local inconsistency from two successive generated frames of Front-Left (FL) camera. Even though our video generation model retains fine 3D consistency, minor discrepancies are inevitable. Our FTGS can effectively reconstruct the scene with awareness of such discrepancy.
4.2 Video Generation with Relative Pose Control